$SPAD/src/input ribbons.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
--the second attempt
drawRibbons(flist, xrange, yrange) ==
sp := createThreeSpace()
num := # flist
yVar := variable yrange
y0:Float    := lo segment yrange
width:Float := (hi segment yrange - y0)/num
for f in flist for color in 1..num repeat
 makeObject(f, xrange, yVar = y0..y0+width,
  var2Steps == 1, colorFunction == (x,y) +-> color,
   space ==sp)
 y0:= y0 + width
vp := makeViewport3D(sp, "Ribbons")
drawStyle(vp, "shade")
outlineRender(vp, "on")
showRegion(vp, "on")
vp

———-

References^{[1] nothing}

4